OpenStudy (anonymous):

Simplify. 6 ----- (sqrt3) + 2

OpenStudy (anonymous):

\[6/\sqrt{3} +2\]

OpenStudy (saifoo.khan):

\[\frac{6}{\sqrt 3 + 2} \times \frac {\sqrt 3 - 2}{\sqrt 3 - 2}\]

OpenStudy (anonymous):

I dont get it. :/

OpenStudy (anonymous):

i agree with saifoo>>>>> taking conjugate

OpenStudy (saifoo.khan):

Whyy? @Algebra1Flunk

OpenStudy (anonymous):

I see it now. 6 / (√3+2) = 6 (√3-2) / (√3+2)(√3-2) = (6√3 - 12) / (√3*√3 - 2*2) = (6√3 - 12) / (3 - 4) = (6√3 - 12) / -1 = 12 - 6√3